OpenStudy (adelyn_helps):

I will fan you if you get the answer right What is the correct solution to the equation? 8(x + 2) = 24 A. 1 B. 3 C. 5 D. 64

OpenStudy (anonymous):

the first step is the following divide both sides by 8 8/8(x+2)=24/8 what u will get?

OpenStudy (adelyn_helps):

I dont know that is why I am asking you

OpenStudy (anonymous):

8/8=?

OpenStudy (adelyn_helps):

1 whole

OpenStudy (anonymous):

yes. and 24/8=?

OpenStudy (anonymous):

8*3 = 24 x+2 = 3

OpenStudy (ninjaslice):

1/1 = 1 whole anything that is like 8/8 will equal 1 whole

OpenStudy (anonymous):

@adelyn_helps what will be 24/8? once you get this one we will countinue with your equation

OpenStudy (adelyn_helps):

3

OpenStudy (anonymous):

yes. so the equation will look like (x+2)=3

OpenStudy (adelyn_helps):

So what would the correct answer be

OpenStudy (anonymous):

now subtract 2 from both sides x+2-2=3-2 1) 2-2 = ? 2) 3-2 =?

OpenStudy (anonymous):

It should be 1 then easy, all you have to do is find out what the x is.

OpenStudy (adelyn_helps):

0 and 1

OpenStudy (anonymous):

yes. x+0=1 or x=1 ..... that's your answer

OpenStudy (anonymous):

8 x 3=24

OpenStudy (adelyn_helps):

@MLG360GABEN your confusing me

OpenStudy (anonymous):

1+2=3, 3 x 8= 24

OpenStudy (ninjaslice):

it is 24. I promise. so the answer is A

OpenStudy (adelyn_helps):

I medaled and fanned you @ksanka

OpenStudy (anonymous):

The answer is A

OpenStudy (ninjaslice):

Due to since there is parentheses yyou have to multiply whatever is in the parentheses. So you need to do 8*x and 8*2 Which then equals 24 when those 2 sums are added. so for 8*1 and 8*2 added together that is 24. hope this helps.

OpenStudy (adelyn_helps):

other person makes more sense